# Hybrid A* Path Planning with Multi-Modal Motion Extension for Four-Wheel Steering Mobile Robots

### v1

@@ Abstract
00footnotetext: * This paper will be submitted to Chinese Control Conference 2026.00footnotetext: * Corresponding authors: Shoukun Wang (bitwsk@bit.edu.cn).00footnotetext: * This work was supported by the National Natural Science Foundation of China under Grant 62473044.
@@ 2.1 Global Path Planning Problem Formulation
Lateral steering employs the same geometric constraint principles as Ackermann steering, but with wheel orientations adjusted perpendicular to the robot’s longitudinal axis, thereby enabling lateral motion. In parallel movement mode, all wheels maintain identical steering angles, allowing the robot to move in any direction without altering its orientation. When the four wheels maintain the same velocity, for the control input angle $\eta$, the motion constraints are:
$\begin{cases}\eta_{1}=\eta_{3}=\eta\\ \eta_{2}=\eta_{4}=\text{sgn}(\eta)(\pi-|\eta|),\end{cases}$ (6)
where $\text{sgn}(\cdot)$ denotes the sign function.
@@ 3.2 Multi-Modal State Representation
The core idea is to extend the original three-dimensional state space $(x,y,\theta)$ to a four-dimensional state space $(x,y,\theta,m)$, where the motion mode variable $m$ is explicitly modeled as the fourth dimension of the state. Specifically, $m\in\{1,2,3\}$ corresponds to Ackermann steering, lateral steering and parallel movement, respectively. Constraint formulas are given in Eqs. 5 and 6.
In the expanded four-dimensional state space, state expansion process includes two operations: intra-modal expansion and inter-modal expansion. Intra-modal expansion maintains the current motion mode unchanged and generates successor states in the three-dimensional pose space according to the corresponding kinematic model; inter-modal expansion performs mode switching at the current position, producing successor states with different motion modes but the same pose. To effectively manage state space complexity, the algorithm employs heuristic modal selection strategies, choosing the most effective motion mode based on the relative direction between the current position and target, and minimizing unnecessary modal switches while maintaining trajectory continuity under constraint satisfaction.
@@ 3.3 Multi-Modal Reeds-Shepp Curves
where $\dot{\phi}(t)$ is the rate of change of wheel angle and $v(t)$ is the reference velocity. This expression indicates that, unlike Ackermann mode where curvature is determined solely by geometric parameters, curvature in parallel mode is directly coupled with velocity: the same wheel angle change rate produces larger curvature at low speeds and corresponds to gentler trajectories at high speeds. This makes parallel movement mode a special case in RS curve extension. Since the robot’s orientation remains constant in this mode, it is only activated when the current orientation aligns with the target orientation or falls within a specified tolerance.
@@ 4.1 Experimental Setup
Max Steer Angle ($\delta_{\max}$) 30° Max Steer Rate ($\dot{\delta}_{\max}$) 180°/s

### v2

@@ Abstract
00footnotetext: * Corresponding authors: Shoukun Wang (bitwsk@bit.edu.cn).00footnotetext: * This work was supported by the National Natural Science Foundation of China under Grant 62473044.
@@ 2.1 Global Path Planning Problem Formulation
Lateral steering preserves the same same geometric constraint principles as the Ackermann mode, but the wheels are reoriented to be perpendicular to the vehicle’s longitudinal axis, allowing the robot to generate pure sideways motion. When the lateral direction is regarded as the effective forward direction, the underlying steering-angle computation becomes structurally identical to that of the Ackermann formulation, with $L$ and $W$ exchanging roles in the geometric constraint. In parallel movement mode, all wheels maintain identical steering angles and velocities, so that the vehicle undergoes pure translation in an arbitrary direction without changing its orientation.
@@ 3.2 Multi-Modal State Representation
The core idea is to extend the original three-dimensional state space $(x,y,\theta)$ to a four-dimensional state space $(x,y,\theta,m)$, where the motion mode variable $m$ is explicitly modeled as the fourth dimension of the state. Specifically, $m\in\{1,2,3\}$ corresponds to Ackermann steering, lateral steering and parallel movement, respectively.
In the expanded four-dimensional state space, state expansion process includes two operations: intra-modal expansion and inter-modal expansion. Intra-modal expansion maintains the current motion mode unchanged and generates successor states in the three-dimensional pose space according to the corresponding kinematic model; inter-modal expansion performs mode switching at the current position, producing successor states with different motion modes but the same pose. To effectively manage state space complexity, the algorithm employs heuristic modal selection strategies, choosing the most effective motion mode based on the relative direction between the current position and target, and minimizing unnecessary modal switches while maintaining trajectory continuity under constraint satisfaction. The sampling strategies associated with each motion mode are illustrated in Fig. 2.
@@ 3.3 Multi-Modal Reeds-Shepp Curves
where $\dot{\phi}(t)$ is the rate of change of wheel angle and $v(t)$ is the reference velocity. This expression indicates that, unlike Ackermann mode where curvature is determined solely by geometric parameters, curvature in parallel mode is directly coupled with velocity: the same wheel angle change rate produces larger curvature at low speeds and corresponds to gentler trajectories at high speeds. This makes parallel movement mode a special case in RS curve extension. Since the robot’s orientation remains constant in this mode, it is only activated when the current orientation aligns with the target orientation or falls within a specified tolerance. The Reeds–Shepp curve visualizations for all motion modes are shown in Fig. 3.
@@ 4.1 Experimental Setup
Max Steer Angle ($\delta_{\max}$) 30° Max Steer Rate ($\dot{\delta}_{\max}$) 360°/s